# Geometric Median Matching for Robust $k$-Subset Selection from Noisy Data

### v1

@@ 1.1 Robustness vs Diversity :
However, in the realistic noisy scenarios involving outliers, this strategy often fails since the noisy examples are wrongly deemed informative for training (Zhang & Sabuncu, 2018; Park et al., 2024). To mitigate this issue, existing pruning methods, tailored for such noisy scenarios, aim to retain the most prototypical / easy samples (Pleiss et al., 2020; Jiang et al., 2018; Har-Peled et al., 2006; Shah et al., 2020; Shen & Sanghavi, 2019). Yet, by prioritizing samples far from the decision boundary, these methods overlook less prototypical but uncorrupted and highly informative examples. This bias introduces a fundamental robustness vs. diversity trade-off (Xia et al., 2022), where emphasizing robustness can lead to a lack of diversity in the pruned subset. This limitation not only results in suboptimal generalization performance but, in extreme cases, can also lead to degenerate solutions (Sugiyama & Kawanabe, 2012).
@@ 2 Related Work
One set of data pruning approaches rely on some carefully designed pruning metrics to rank the training samples based on the scores and retain a fraction of them as representative samples, used for training the downstream model. For example, (Xia et al., 2022; Joshi & Mirzasoleiman, 2023; Sorscher et al., 2022) calculate the importance score of a sample in terms of the distance from the centroid of its corresponding class marginal. Samples closer to the centroid are considered most prototypical (easy) and those far from the centroid are treated as least prototypical (hard).
While this work primarily focuses on spatial approaches, it is worth mentioning that the canonical importance scoring criterion have been proposed in terms gradient norm (Paul et al., 2021; Needell et al., 2014), uncertainty (Pleiss et al., 2020) and forgetfulness (Toneva et al., 2018; Feldman & Zhang, 2020). Typically, samples closer to the class centroid in feature space tend to have lower gradient norms, exhibit lower uncertainty, and are harder to forget during training. In contrast, samples farther from the centroid generally have higher gradient norms, greater uncertainty, and are easier to forget (Paul et al., 2021; Sorscher et al., 2022; Xia et al., 2022).
@@
Another paradigm selects subsets by explicitly matching certain statistical moments or characteristics of the full dataset. Rather than selecting examples purely by rank, these methods formulate subset selection as an optimization problem: find a (possibly weighted) subset such that some property (loss, gradient, feature distribution) of the subset approximates that of the entire data.
A second set of works re-formulate this problem as minimizing a moment matching objective (Chen et al., 2010; Campbell & Broderick, 2018; Dwivedi & Mackey, 2021) that aims to select a subset whose mean closely matches that of the entire dataset.
Moreover, (Mirzasoleiman et al., 2020) extended the moment-matching approach to the gradient space, selecting subsets that preserve the overall gradient statistics of the full dataset.
@@
While there has been considerable progress in designing subset selection algorithms for large-scale, clean datasets, extending these methods to noisy or corrupted settings remains relatively under-explored. As we show in Lemma 3, classical approaches remain vulnerable under the arbitrary corruption regime. Since the empirical mean is not robust in adversarial or heavy-tailed settings, any subset selection method relying on it inherits the same vulnerability.
On the other hand, most existing robust methods are only designed for specific noise types; for instance, (Park et al., 2024; Pleiss et al., 2020) tackles the problem for label noise. However, the specialized approach of re-labeling does not extend to arbitrary noise settings.
Most existing solutions mitigate noise by selecting the most prototypical or easy examples (Shah et al., 2020), effectively discarding noisy samples. However, as mentioned earlier, these methods lack diversity by ignoring boundary examples, which can be crucial for decision-making.
More recently, (Xia et al., 2022) proposed an alternative strategy: instead of selecting only the easiest or hardest samples, they compute the distance of each sample from the class centroid, sort the scores, and retain those closest to the median score. While this approach introduces a form of moderation, it still fundamentally relies on distance from the empirical mean.
To the best of our knowledge, we are the first to provide theoretical robustness guarantees for subset selection under gross corruption settings, moving beyond prior heuristic methods. While previous works (Thompson, 2022; Xu et al., 2025) have explored robustness, they are limited to restrictive theoretical settings and linear models. Our approach generalizes to broader, real-world applications.
@@
Kernel Herding (Welling, 2009; Welling & Chen, 2010; Chen & Welling, 2010; Chen et al., 2010; Bach et al., 2012) is an approach originally developed to produce a small set of “representative” points that approximate the mean embedding of a probability distribution within a reproducing kernel Hilbert space (RKHS). Specifically, if $\mu$ denotes the true mean embedding of the data distribution $\mathcal{D}$, Kernel Herding iteratively constructs a $k$-subset of data points $\{x_{i}\}$ whose empirical kernel mean $\hat{\mu}_{k}$ converges to $\mu$ in RKHS norm at a rate of $\mathcal{O}(1/k)$. This deterministic sampling approach can be viewed as an alternative to standard Monte Carlo or MCMC sampling, often yielding lower-variance estimates for certain integrals.
The proposed GmMatching method shares the same moment-matching spirit as Kernel Herding—namely, creating a subset whose “mean” (in some appropriately defined sense) is close to that of the full dataset. However, the crucial difference lies in which moment is being matched. While Kernel Herding aims to match the empirical mean of the original dataset (potentially sensitive to outliers or heavy-tailed distributions), GmMatching targets a robust moment Gm. By substituting the empirical mean with a robust moment estimator, GmMatching inherits the same desirable convergence rate of $\mathcal{O}(1/k)$ from Kernel Herding, while becoming significantly less sensitive to outliers or contaminated data. Intuitively, GmMatching leverages the same sequential greedy selection process used in Kernel Herding, but bases its updates on robust statistics—thus ensuring that a small subset can still capture the “core” of the distribution even under less ideal data conditions.
Because GmMatching inherits the theoretical machinery of Kernel Herding (but replaces the empirical mean with Gm), it enjoys the same $\mathcal{O}(1/k)$ rate while improving resilience to noise and outliers. This robust formulation is especially advantageous for large-scale real-world datasets that may include mislabeled examples or exhibit skewed distributions.
@@ 4.1 Vulnerability under Gross Corruption
5 Geometric Median (Gm) Matching
@@ Appendix D Proof of Theorem 1
D.1 Bounding Estimation Error from Gm

### v2

@@ 1.1 Robustness vs Diversity :
However, in the realistic noisy scenarios involving outliers, this strategy often fails since the noisy examples are wrongly deemed informative for training (Zhang & Sabuncu, 2018; Park et al., 2024). To mitigate this issue, existing pruning methods, tailored for such noisy scenarios, aim to retain the most prototypical / easy samples (Pleiss et al., 2020; Jiang et al., 2018; Har-Peled et al., 2007; Shah et al., 2020; Shen & Sanghavi, 2019). Yet, by prioritizing samples far from the decision boundary, these methods overlook less prototypical but uncorrupted and highly informative examples. This bias introduces a fundamental robustness vs. diversity trade-off (Xia et al., 2022), where emphasizing robustness can lead to a lack of diversity in the pruned subset. This limitation not only results in suboptimal generalization performance but, in extreme cases, can also lead to degenerate solutions (Sugiyama & Kawanabe, 2012).
@@ 2 Related Work
One set of data pruning approaches rely on some carefully designed pruning metrics to rank the training samples based on the scores and retain a fraction of them as representative samples, used for training the downstream model. For example, (Xia et al., 2022; Joshi & Mirzasoleiman, 2023; Sorscher et al., 2022) calculate the importance score of a sample in terms of the distance from the centroid of its corresponding class marginal. Samples closer to the centroid are considered most prototypical (easy) and those far from the centroid are treated as least prototypical (hard). While this work primarily focuses on spatial approaches, it is worth mentioning that the canonical importance scoring criterion have been proposed in terms gradient norm (Paul et al., 2021; Needell et al., 2014), uncertainty (Pleiss et al., 2020) and forgetfulness (Toneva et al., 2018; Feldman & Zhang, 2020). Typically, samples closer to the class centroid in feature space tend to have lower gradient norms, exhibit lower uncertainty, and are harder to forget during training. In contrast, samples farther from the centroid generally have higher gradient norms, greater uncertainty, and are easier to forget (Paul et al., 2021; Sorscher et al., 2022; Xia et al., 2022). However, such scoring-based selection methods typically rely on the empirical mean to define centroids, making them brittle under data corruption—a vulnerability that becomes critical in noisy or adversarial settings.
@@
A second family of methods tackles data selection as an optimization problem, aiming to match certain statistical moments—such as means, losses, or gradients—between the subset and the full dataset. These techniques, often rooted in coreset construction, select a subset whose empirical properties approximate those of the entire dataset (Chen et al., 2010; Campbell & Broderick, 2018; Dwivedi & Mackey, 2021). A notable extension is gradient-based moment matching (Mirzasoleiman et al., 2020), which seeks to preserve the full dataset’s gradient dynamics within the selected subset, enabling efficient yet faithful training.
@@
Despite substantial advances in subset selection algorithms for large-scale datasets, most existing methods are designed for idealized, clean training data. In practical scenarios, however, real-world datasets are often noisy—affected by mislabeled examples, corrupted features, or adversarial attacks. Unfortunately, extending traditional subset selection methods to such noisy settings remains an under-explored challenge. A key limitation of many classical approaches lies in their dependence on the empirical mean—a statistic known to be highly sensitive to outliers and heavy-tailed distributions. As we formally demonstrate in Lemma 3, this renders them brittle under arbitrary corruption, where even a small fraction of adversarial points can severely skew the selection process. Consequently, any method that uses the empirical mean as a reference—either directly (e.g., centroid-based selection) or indirectly (e.g., gradient- or uncertainty-based ranking)—inherits the same fundamental vulnerability.
Several recent works have attempted to address robustness, but often in narrowly scoped settings. For instance, methods tailored to label noise (Pleiss et al., 2020; Park et al., 2024) typically rely on techniques such as relabeling or sample reweighting, assuming access to a relatively clean subset or side information. These strategies can be effective in controlled scenarios but do not generalize well to broader corruption types such as input perturbations, distribution shift, or structural outliers.
Another class of methods mitigate noise by aggressively pruning away uncertain or atypical examples—often selecting only the most prototypical or easy-to-learn samples (Shah et al., 2020; Toneva et al., 2018; Jiang et al., 2018; Har-Peled et al., 2007). While such selection improves robustness by avoiding noisy outliers, it inadvertently sacrifices diversity, discarding informative hard examples that are essential for learning decision boundaries. This introduces a well-known robustness–diversity tradeoff (Xia et al., 2022; Feldman & Zhang, 2020), where safe selection leads to suboptimal generalization. To balance this trade-off, Xia et al. (2022) proposed a moderation strategy that retains samples closest to the median distance from the class centroid—aiming to avoid both overly hard and overly easy examples. However, this approach still fundamentally relies on spatial distances measured from the empirical mean, making it vulnerable to corruption that distorts this reference point. Furthermore, while a few recent works have studied robust subset selection in a theoretical manner, they are either limited to linear models(Xu et al., 2025) or assume simplified theoretical frameworks(Thompson, 2022; Park et al., 2024), strong assumptions (Qian et al., 2017). These works stop short of providing practical algorithms or generalization guarantees applicable to modern, highly nonlinear deep learning systems.
In contrast, our work introduces a simple, theoretically grounded algorithm for subset selection under arbitrary data corruption, bridging the gap between robustness and diversity. GmMatching generalizes to deep models and real-world noisy data, offering provable guarantees under the Gross Corruption.
@@
GmMatching builds upon the classical Kernel Herding framework (Welling, 2009; Chen & Welling, 2010; Bach et al., 2012), which selects representative subsets that match the mean embedding of the data distribution in a reproducing kernel Hilbert space (RKHS). While Kernel Herding offers favorable convergence guarantees ${\mathcal{O}}(1/k)$ vs. ${\mathcal{O}}(1/\sqrt{k})$ for random sampling; it relies on the empirical mean, making it vulnerable to outliers and adversarial corruption. GmMatching preserves the moment-matching spirit of Kernel Herding but replaces the empirical mean with the geometric median (GM), a robust estimator with a breakdown point of 1/2. This simple substitution significantly improves robustness without compromising convergence. To our knowledge, GmMatching is the first method to combine robust estimation with herding-style greedy selection, enabling provably fast and stable subset selection even under high rates of data corruption.
@@ 4.1 Vulnerability under Gross Corruption
5 Geometric Median Matching
@@ Appendix D Proof of Theorem 1
D.1 Bounding Estimation Error from Approximate Geometric Median